# Exploring Ground and Excited States via Single Reference Coupled-Cluster Theory and Algebraic Geometry

### v1

@@ Abstract
The exploration of the root structure of coupled cluster equations holds both foundational and practical significance for computational quantum chemistry. This study provides insight into the intricate root structures of these non-linear equations at both the CCD and CCSD level of theory. We utilize computational techniques from algebraic geometry, specifically the monodromy and parametric homotopy continuation methods, to calculate the full solution set. We compare the computed CC roots against various established theoretical upper bounds, shedding light on the accuracy and efficiency of these bounds. We hereby focus on the dissociation processes of four-electron systems such as (H2)2 in both D2h and D∞h configurations, H4 symmetrically distorted on a circle, and lithium hydride. We moreover investigate the ability of single-reference coupled cluster solutions to approximate excited state energies. We find that multiple CC roots describe energies of excited states with high accuracy. Remarkably, our investigations reveal that for systems like lithium hydride, CC not only provides high-accuracy approximations to several excited state energies but also to the states themselves.
@@ 1 Introduction
Especially the last question is controversial, to say the least. On one side, it can be proven that as long a state is not strictly orthogonal to the chosen reference state, CC theory can – in its untruncated form – describe that state exactly. On the other side, excited states commonly have poor overlap with the Hartree-Fock ground state reference – hindering the convergence of (quasi) Newton-type methods. Besides such theoretical arguments, it has been numerically shown that higher-order CC solutions can approximate excited state energies well while the CC solutions provide poor approximations to the actual states 13. Conventionally, excited states are computed in a post-single-reference CC manner. This is done by either constructing and diagonalizing the similarity-transformed Hamiltonian in equation-of-motion CC 14, 15, 16, 17, 18, 19, 20 or by examining the poles of the linear-response function in linear-response CC 16, 21, 22, 23. In either case, the excited states provided by these methods are inherently biased toward the lowest eigenstate upon which they are built. As a direct consequence, the accuracy of such approaches is known to be quite poor for certain classes of eigenstates that are energetically higher than the ground state 24, 25, 26, 27, 28, 29.
Besides the fundamental nature of questions i)–iii), the existence of multiple roots to the CC equations can present challenges for the practical implementations of the method. For example, the roots may be difficult to adequately converge to 5, 30, 31, 32, and the convergence properties of the iterative solution methods can strongly depend on the employed initial guess 33, 34. See also Ref. 35 for a mathematical exposition of this problem and Ref. 36 for a potential remedy.
@@ 1.1 Previous works and perspective
We highlight that the inaugural investigation of the root structure of CC methods dates back to 1978 when Zivkovic and Monkhorst scrutinized the singularities and multifarious solutions in single-reference CC equations 37. Building upon this work, Paldus and colleagues subsequently conducted mathematical and numerical analyses during the early 1990s, focusing on the manifold of solutions of both single-reference and state-universal multi-reference CC equations, while also elucidating their singularities and analytical characteristics 38, 39. In 1998, Kowalski and Jankowski revived homotopy methods in connection with the single-reference CC theory and used them to solve a CCD system 40. Kowalski with Piecuch then extended the application of homotopy methods to the CCSD, CCSDT, and CCSDTQ 41 for a 4-electron system described in a minimal basis set. From these investigations, they derived the $\beta$-nested equations and proved the Fundamental Theorem of the $\beta$-NE Formalism 42, which enabled an explanation of curves connecting multiple solutions of the various CC polynomial systems, i.e. from CCSD to CCSDT, CCSDT to CCSDTQ, etc. This Kowalski-Piecuch homotopy was recently analyzed in-depth using topological degree theory 43, 44. Moreover, Kowalski with Piecuch used homotopy methods to obtain all solutions of the generalized Bloch equation 45 and explain symmetry breaking processes in the CC formulation 46, 47, 48, 49, 50. It is noteworthy that these fundamental results of CC theory have spurred the development of the method of moment CC approach 51, 52, 53, 54, which in turn has led to the highly efficient completely renormalized CC approach 55, 56, 57. This emphasizes the pivotal role of this research trajectory in driving methodological innovations within CC theory. A perspective on the recent use of homotopy methods in the context of quantum chemistry can be found in Ref. 58.
@@
A more algebraic and mathematically driven approach to this problem has recently gained momentum. Initialized by some of the authors, a first algebraic perspective was introduced in Ref. 59. This was followed by an in-depth algebraic investigation leading to the introduction of the CC truncation varieties 60 – which are the subject of this work. This work moreover contained open mathematical questions and conjectures, one of which was proven shortly after 61. More generally, these works are part of a presently increasing interest in the mathematical understanding of CC theory, driven by the applied mathematics community. In this context, Schneider performed the first local analysis of CC theory in 2009 employing a functional analytic framework 62. Subsequently, this approach was refined 63, 64 and applied to different CC methods 65, 66, 67. A more versatile framework for analyzing general CC variants using topological degree theory was later proposed by Csirik and Laestadius 43, 44. The most recent numerical analysis results regarding single reference CC were established by Hassan, Maday, and Wang 68, 69.
@@ 2 Coupled-cluster theory
The cluster operator is a linear combination of elementary particle-hole excitation operators, i.e.
@@
where we highlight the dependence of $\hat{T}$ on its expansion coefficients by writing $\hat{T}(\bf t)$. Recall that elementary particle-hole excitation operators $\hat{X}_{\mu}$ are merely compositions of projections onto a subset of single-particle basis elements (the occupied orbitals) and wedge products with another subset of single-particle basis elements (the virtual orbital). It is therefore convenient to label the elementary particle-hole excitation operators using multi-indices $\mu>0$ that clarify the projections and wedge products involved 12, 62. Therefore, acting on the reference state $|{\Phi_{0}}\rangle$, the elementary particle-hole excitation operators define the Hilbert space of functions that are $L^{2}$-orthogonal to $|{\Phi_{0}}\rangle$. We henceforth set $\mathcal{V}:=\{|{\Phi_{0}}\rangle\}^{\perp}$ and note that $\hat{T}({\bf t})|{\Phi_{0}}\rangle\in\mathcal{V}$ for all CC amplitudes $\mathbf{t}$. Hence, we can express Eq. (4) in the more common form as a set of projective equations, i.e.
@@
We conclude this section by noting that in the case of untruncated CC, amplitudes that solve the full CC equations describe eigenstates of the Hamiltonian. Therefore, the cardinality of $\mathcal{S}$ is exactly the number of eigenstates of $\hat{H}$ that are intermediately normalized and the number of roots to the CC equations is bounded from above by the number of Slater determinants. However, when truncations are imposed, the cardinality of $\mathcal{S}$ increases, i.e. truncated CC theory yields (some) unphysical solutions. Therefore, it becomes less clear what the different elements in $\mathcal{S}$ describe. Understanding the variety $\mathcal{S}$ and characterizing (some of) its elements are the subject of this manuscript.
@@ 3 Truncation varieties
In CC theory one encounters various algebraic equations that define different varieties. The truncation varieties 60 arise from the exponential parametrization, i.e.
${\rm exp}:\mathbb{V}\to\mathcal{V},~{}\mathbf{t}\mapsto e^{\hat{T}(\mathbf{t})}|\Phi_{0}\rangle,$ (9)
where $\mathbb{V}$ is the CC amplitude space. In the untruncated CC method, the exponential map is bijective. However, when truncations are introduced, this property no longer holds. Truncations correspond to restrictions of the exponential map to a subspace, e.g. for CCSD the exponential map is restricted to $\mathbb{V}_{\{1,2\}}\subset\mathbb{V}$ consisting of cluster amplitudes with only one and two particle-hole excitations. The corresponding truncation variety, $V_{\{1,2\}}$ is then defined as the smallest variety containing ${\rm exp}(\mathbb{V}_{\{1,2\}})$. This construction of the truncation variety can be generalized to any CC truncation: Let $\sigma$ denote the CC truncation rank, i.e. $\sigma\subset[\![N]\!]=\{1,\dots,N\}$, where $N$ denotes the number of electrons in the system. Then, the smallest variety containing ${\rm exp}(\mathbb{V}_{\sigma})$ is the $\sigma$-truncation variety denoted $V_{\sigma}$.
Remarkably, the truncation varieties can be used to express the projective CC equations (6) as an eigenvalue problem, i.e.
@@
This formulation is indeed equivalent to the projective CC equations in Eq. (6) for CCD, as well as for rank complete CC truncations, i.e. CCSD, CCSDT, CCSDTQ etc., see Theorem 5.11 in Ref. 60. The number of roots for these CC variants is bounded by the number of solutions to Eq. (10) for a generic Hamiltonian matrix $\hat{H}$. We denote this number the CC degree of the corresponding truncation variety. The CC degree serves as a complexity measure for solving the CC equations and Bézout’s theorem provides the upper bound
@@
where $\dim(V_{\sigma})$ is the dimension of the truncation variety and $\deg(V_{\sigma})$ is its degree. The dimension of the truncation variety is well understood. The reason is that cluster analysis allows us to bijectively map cluster amplitudes to CI-coefficients using polynomial equations, see e.g. Refs. 12, 70, 32 or Proposition 2.4 in Ref. 60 for a mathematical proof. Then, since $V_{\sigma}$ arises as a deformation of the linear space $\mathbb{V}_{\sigma}$ under a polynomial with polynomial inverse, the dimension of $V_{\sigma}$ is equal to the dimension of $\mathbb{V}_{\sigma}$. For example in the case of CCSD, we obtain
@@
The degree of ${V}_{\sigma}$ on the other hand is not as easily computable and requires a more profound understanding of ${V}_{\sigma}$. For special cases, however, explicit formulas for the degree have been established; for example, for CCS, the truncation variety is the Grassmanian (see Theorem 3.5 in Ref. 60) which is mathematically well understood. In other cases we rely either on symbolic methods, provided by e.g. Macaulay2 71, or on numerical methods, accessible in e.g. HomotopyContinuation.jl 72, to compute the degree.
Despite the complexity of calculating $\deg(V_{\sigma})$ the bound given by Eq. (12) provides significant insight into bounding the CC degree, since it is less complex to calculate $\deg(V_{\sigma})$ than to calculate the CC degree directly. Moreover, this upper bound is notably better than the previously known upper bounds obtained by applying Bézout’s theorem directly to the projective CC equations, see Section 6 in Ref. 60.
@@ 4 Computational methods
$\mathcal{F}_{\rm CC}(\sigma)=\{[(\hat{G}-E\cdot I_{K})e^{\hat{T}(\mathbf{t})}]_{\sigma}~{}\big{|}~{}\hat{G}\in\mathbb{F}^{K\times K}\},$ (14)
where the energy $E\in\mathbb{C}$ and cluster amplitudes $\mathbf{t}\in\mathbb{V}_{\sigma}$ are the variables, and the entries of the matrix $\hat{G}$ are the parameters. We will ultimately set $\hat{G}=\hat{H}$ to be the Hamiltonian of interest, but $\mathcal{F}_{\rm CC}(\sigma)$ is a much broader family of CC-like systems with $\hat{G}\in\mathbb{F}^{K\times K}$ an arbitrary matrix. The parameter continuation theorem provides an upper bound to the number of roots of any polynomial system $F_{\hat{G}}\in\mathcal{F}_{\rm CC}$. Moreover, this bound is tight and equality is obtained for all polynomial systems with parameters in a certain dense subspace of $\mathbb{F}^{K\times K}$. We subsequently assume the validity of the parameter continuation theorem and refer the interested reader to Section 3 in Ref. 73. The upper bound provided by the parameter continuation theorem yields the CC degree mentioned in the previous section.
@@
We emphasize that, at their core, both methods use a homotopy to track new solutions from known solutions 74, 75, 76, 77. More precisely, we define a homotopy for a (piecewise smooth) path $\gamma:[0,1]\to\mathbb{F}^{K\times K}$ in the parameter space via
@@
where $H(E,\mathbf{t},\lambda)$ describes the continuous deformation of a start system $G(E,\mathbf{t})=H(E,\mathbf{t},1)$ to a target system $F(E,\mathbf{t})=H(E,\mathbf{t},0)$. The individual solution paths to the system $H(E,\mathbf{t},\lambda)$ are then tracked from $G(E,\mathbf{t})$ to $F(E,\mathbf{t})$ as $\lambda\to 0$. This is accomplished via an ordinary differential equation, known as the Davidenko differential equation 78, 79, i.e.
@@ 4.1 Monodromy
This approach can be effectively utilized to find all solutions to a generic system of equations. For instance, consider a scenario where the only known solution for the polynomial in Eq. (18) is $x_{2}=2$. By tracking this solution along the unit circle, we can determine another solution, $x_{1}=1$. When we extend this tracking across sufficiently many loops, the method reveals all possible roots of the system.
This process is central to the monodromy solver. The solver begins with a single known solution and numerically traces known solutions through multiple loops within the parameter space. This tracking continues until no additional solutions are found over a predetermined number of loops, leading to the conclusion that all possible solutions to the system have been numerically identified.

### v2

@@ Abstract
The exploration of the root structure of coupled cluster equations holds both foundational and practical significance for computational quantum chemistry. This study provides insight into the intricate root structures of these non-linear equations at both the CCD and CCSD level of theory. We utilize computational techniques from algebraic geometry, specifically the monodromy and parametric homotopy continuation methods, to calculate the full solution set. We compare the computed CC roots against various established theoretical upper bounds, shedding light on the accuracy and efficiency of these bounds. We hereby focus on the dissociation processes of four-electron systems such as (H2)2 in both D2h and D∞h configurations, H4 symmetrically distorted on a circle, and lithium hydride. We moreover investigate the ability of single-reference coupled cluster solutions to approximate excited state energies. We find that multiple CC roots describe energies of excited states with high accuracy. Our investigations reveal that for systems like lithium hydride, CC not only provides high-accuracy approximations to several excited state energies but also to the states themselves.
[…]
supporting_information
@@ 1 Introduction
Especially the last question is controversial, to say the least. On one side, it can be proven that as long a state is not strictly orthogonal to the chosen reference state, CC theory can – in its untruncated form – describe that state exactly. On the other side, excited states commonly have poor overlap with the Hartree-Fock ground state reference – hindering the convergence of (quasi) Newton-type methods. Besides such theoretical arguments, it has been numerically shown that higher-order CC solutions can approximate excited state energies well while the CC solutions provide poor approximations to the actual states 13. Conventionally, excited states are computed in a post-single-reference CC manner. This is done by either constructing and diagonalizing the similarity-transformed Hamiltonian in equation-of-motion CC 14, 15, 16, 17, 18, 19, 20, or by examining the poles of the linear-response function in linear-response CC 16, 21, 22, 23. In either case, the excited states provided by these methods are inherently biased toward the lowest eigenstate upon which they are built. As a direct consequence, the accuracy of such approaches is known to be quite poor for certain classes of eigenstates that are energetically higher than the ground state 24, 25, 26, 27, 28, 29. An alternative approach, which aligns more closely with the ideas discussed in this work and offers greater accuracy for doubly excited states compared to EOM-CC30, is known as $\Delta$CC. This method directly employs the CC wavefunction parametrization for excited states31, 32, 33, 34, 35, 36, 37, 38, 39, 40.
Besides the fundamental nature of questions i)–iii), the existence of multiple roots to the CC equations can present challenges for the practical implementations of the method. For example, the roots may be difficult to adequately converge to 5, 41, 42, 43, 30, and the convergence properties of the iterative solution methods can strongly depend on the employed initial guess 44, 45. See also Ref. 46 for a mathematical exposition of this problem and Ref. 47 for a potential remedy.
@@ 1.1 Previous works and perspective
We highlight that the inaugural investigation of the root structure of CC methods dates back to 1978 when Zivkovic and Monkhorst scrutinized the singularities and multifarious solutions in single-reference CC equations 48. Building upon this work, Paldus and colleagues subsequently conducted mathematical and numerical analyses during the early 1990s, focusing on the manifold of solutions of both single-reference and state-universal multi-reference CC equations, while also elucidating their singularities and analytical characteristics 49, 50. In 1998, Kowalski and Jankowski revived homotopy methods in connection with the single-reference CC theory and used them to solve a CCD system 51. Kowalski with Piecuch then extended the application of homotopy methods to the CCSD, CCSDT, and CCSDTQ 13 for a 4-electron system described in a minimal basis set. From these investigations, they derived the $\beta$-nested equations and proved the Fundamental Theorem of the $\beta$-NE Formalism 13, which enabled an explanation of curves connecting multiple solutions of the various CC polynomial systems, i.e. from CCSD to CCSDT, CCSDT to CCSDTQ, etc. This Kowalski-Piecuch homotopy was recently analyzed in-depth using topological degree theory 52, 53. Moreover, Kowalski with Piecuch used homotopy methods to obtain all solutions of the generalized Bloch equation 54 and to compute all solutions to the multireference coupled-cluster equations within the framework of the state-universal coupled-cluster formalism55. These works extend the preceding works by Kowalski and Jankowski56, 35, 36, 37, 57. It is noteworthy that these fundamental results of CC theory have spurred the development of the method of moment CC approach 58, 59, 60, 61, which in turn has led to the highly efficient completely renormalized CC approach 62, 63, 64. This emphasizes the pivotal role of this research trajectory in driving methodological innovations within CC theory. A perspective on the recent use of homotopy methods in the context of quantum chemistry can be found in Ref. 65.
@@
A more algebraic and mathematically driven approach to this problem has recently gained momentum. Initialized by some of the authors, a first algebraic perspective was introduced in Ref. 66. This was followed by an in-depth algebraic investigation leading to the introduction of the CC truncation varieties 67 – which are the subject of this work. This work moreover contained open mathematical questions and conjectures, one of which was proven shortly after 68. More generally, these works are part of a presently increasing interest in the mathematical understanding of CC theory, driven by the applied mathematics community. In this context, Schneider performed the first local analysis of CC theory in 2009 employing a functional analytic framework 69. Subsequently, this approach was refined 70, 71 and applied to different CC methods 72, 73, 74. A more versatile framework for analyzing general CC variants using topological degree theory was later proposed by Csirik and Laestadius 52, 53. The most recent numerical analysis results regarding single reference CC were established by Hassan, Maday, and Wang 75, 76.
@@ 2 Coupled-cluster theory
The cluster operator is a linear combination of elementary particle-hole excitation operators5, i.e.
@@
where we highlight the dependence of $\hat{T}$ on its expansion coefficients by writing $\hat{T}(\bf t)$. It is therefore convenient to label the elementary particle-hole excitation operators using multi-indices $\mu>0$ that clarify the projections and wedge products involved 12, 69. Therefore, acting on the reference state $|{\Phi_{0}}\rangle$, the elementary particle-hole excitation operators define the Hilbert space of functions that are $L^{2}$-orthogonal to $|{\Phi_{0}}\rangle$. We henceforth set $\mathcal{V}:=\{|{\Phi_{0}}\rangle\}^{\perp}$ and note that $\hat{T}({\bf t})|{\Phi_{0}}\rangle\in\mathcal{V}$ for all CC amplitudes $\mathbf{t}$. Hence, we can express Eq. (4) in the more common form as a set of projective equations, i.e.
@@
We conclude this section by noting that in the case of untruncated CC, amplitudes that solve the full CC equations describe eigenstates of the Hamiltonian. Therefore, the cardinality of $\mathcal{S}$ is exactly the number of eigenstates of $\hat{H}$ that are intermediately normalized and the number of roots to the CC equations is bounded from above by the number of Slater determinants. However, when truncations are imposed, the cardinality of $\mathcal{S}$ increases, and can ultimately lead to more solutions than provided by full diagonalization of the system. Consequently, it becomes less clear what the different elements in $\mathcal{S}$ describe. Understanding the variety $\mathcal{S}$ and characterizing (some of) its elements are the subject of this manuscript.
@@ 3 Truncation varieties
In CC theory one encounters various algebraic equations that define different varieties. The truncation varieties 67 arise from the exponential parametrization, i.e.
${\rm exp}:\mathbb{V}\to\mathcal{H},~{}\mathbf{t}\mapsto e^{\hat{T}(\mathbf{t})}|\Phi_{0}\rangle,$ (9)
where $\mathbb{V}$ is the CC amplitude space and $\mathcal{H}$ denotes the space of intermediately normalized wave functions. In the untruncated CC method, the exponential map is bijective. However, when truncations are introduced, this property no longer holds. Truncations correspond to restrictions of the exponential map to a subspace, e.g. for CCSD the exponential map is restricted to $\mathbb{V}_{\{1,2\}}\subset\mathbb{V}$ consisting of cluster amplitudes with only one and two particle-hole excitations. The corresponding truncation variety, $V_{\{1,2\}}$ is then defined as the smallest variety containing ${\rm exp}(\mathbb{V}_{\{1,2\}})$. This construction of the truncation variety can be generalized to any CC truncation: Let $\sigma$ denote the CC truncation rank, i.e. $\sigma\subset[\![N]\!]=\{1,\dots,N\}$, where $N$ denotes the number of electrons in the system. Then, the smallest variety containing ${\rm exp}(\mathbb{V}_{\sigma})$ is the $\sigma$-truncation variety denoted $V_{\sigma}$.
In this article, we employ the unlinked CC equations , which correspond to
@@
This formulation is equivalent to the projective CC equations in Eq. (6) for CCD, as well as for rank complete CC truncations, i.e. CCSD, CCSDT, CCSDTQ etc., see e.g. Ref. 77 for the original proof of this statement or Theorem 5.11 in Ref. 67. The number of roots for these CC variants is bounded by the number of solutions to Eq. (10) for a generic Hamiltonian matrix $\hat{H}$. We denote this number the CC degree of the corresponding truncation variety. The CC degree serves as a complexity measure for solving the CC equations and Bézout’s theorem provides the upper bound
@@
where $\dim(V_{\sigma})$ is the dimension of the truncation variety and $\deg(V_{\sigma})$ is its degree. The dimension of the truncation variety is well understood. The reason is that cluster analysis allows us to bijectively map cluster amplitudes to CI-coefficients using polynomial equations, see e.g. Refs. 12, 78, 43 or Proposition 2.4 in Ref. 67 for a mathematical proof. Then, since $V_{\sigma}$ arises as a deformation of the linear space $\mathbb{V}_{\sigma}$ under a polynomial with polynomial inverse, the dimension of $V_{\sigma}$ is equal to the dimension of $\mathbb{V}_{\sigma}$. For example in the case of CCSD, we obtain
@@
The degree of ${V}_{\sigma}$ on the other hand is not as easily computable and requires a more profound understanding of ${V}_{\sigma}$. For special cases, however, explicit formulas for the degree have been established; for example, for CCS, the truncation variety is the Grassmanian (see Theorem 3.5 in Ref. 67) which is mathematically well understood. In other cases we rely either on symbolic methods, provided by e.g. Macaulay2 79, or on numerical methods, accessible in e.g. HomotopyContinuation.jl 80, to compute the degree.
Despite the complexity of calculating $\deg(V_{\sigma})$ the bound given by Eq. (12) provides significant insight into bounding the CC degree, since it is less complex to calculate $\deg(V_{\sigma})$ than to calculate the CC degree directly. Moreover, this upper bound is notably better than the previously known upper bounds obtained by applying Bézout’s theorem directly to the projective CC equations, see Section 6 in Ref. 67.
@@ 4 Computational methods
$\mathcal{F}_{\rm CC}(\sigma)=\{[(\hat{G}-E\cdot I_{K})e^{\hat{T}(\mathbf{t})}|\Phi_{0}\rangle]_{\sigma}~{}\big{|}~{}\hat{G}\in\mathbb{F}^{K\times K}\},$ (14)
where the energy $E\in\mathbb{C}$ and cluster amplitudes $\mathbf{t}\in\mathbb{V}_{\sigma}$ are the variables, and the entries of the matrix $\hat{G}$ are the parameters. We will ultimately set $\hat{G}=\hat{H}$ to be the Hamiltonian of interest, but $\mathcal{F}_{\rm CC}(\sigma)$ is a much broader family of CC-like systems with $\hat{G}\in\mathbb{F}^{K\times K}$ an arbitrary matrix. Note that the matrix $\hat{G}$ can be parameterized including further quantum mechanical structures, however, experiments have shown that this does not result in significantly fewer roots while yielding slower calculations due to the inclusions of further constraints. The parameter continuation theorem provides an upper bound to the number of roots of any polynomial system $F_{\hat{G}}\in\mathcal{F}_{\rm CC}$. Moreover, this bound is tight and equality is obtained for all polynomial systems with parameters in a certain dense subspace of $\mathbb{F}^{K\times K}$. We subsequently assume the validity of the parameter continuation theorem and refer the interested reader to Section 3 in Ref. 81. The upper bound provided by the parameter continuation theorem yields the CC degree mentioned in the previous section.
@@
We emphasize that, at their core, both methods use a homotopy to track new solutions from known solutions 82, 83, 84, 85. More precisely, we define a homotopy for a (piecewise smooth) path $\gamma:[0,1]\to\mathbb{F}^{K\times K}$ in the parameter space via
@@
where $H(E,\mathbf{t},\lambda)$ describes the continuous deformation of a start system $G(E,\mathbf{t})=H(E,\mathbf{t},1)$ to a target system $F(E,\mathbf{t})=H(E,\mathbf{t},0)$. The individual solution paths to the system $H(E,\mathbf{t},\lambda)$ are then tracked from $G(E,\mathbf{t})$ to $F(E,\mathbf{t})$ as $\lambda\to 0$. This is accomplished via an ordinary differential equation, known as the Davidenko differential equation 86, 87, i.e.
@@ 4.1 Monodromy
This approach can be effectively utilized to find all solutions to a generic system of equations. For instance, consider a scenario where the only known solution for the polynomial in Eq. (18) is $x_{2}=2$. By tracking this solution along $\gamma$, we can determine another solution, $x_{1}=1$. When we extend this tracking across sufficiently many loops, the method reveals all possible roots of the system.
This process is central to the monodromy solver. The solver begins with a single known solution and numerically tracks the solution through a loop within the parameter space to find a new solution. The new known solution can now be tracked through the same loop. This procedure is continued until an already-known solution is found. Then we generate a new loop and repeat this for all known solutions.
To terminate this procedure a stopping criterion for the monodromy solver has to be chosen. As described above the loops in the parameter space define a permutation of the solution set. Therefore the probability of obtaining all solutions iterating through a fixed number of loops is equal to the probability that a fixed number of random permutations in the symmetry group of the solution set $R=\{\mathbf{x}_{1},\dots,\mathbf{x}_{M}\}$, denoted by $S_{R}$, generate a transitive subgroup. Recall that a transitive subgroup in $S_{R}$ is a subgroup such that we can permute any element in $R$ to another by a permutation in that subgroup. Dixon88 shows the probability that an ordered pair of random permutations in $S_{R}$ generates a transitive subgroup is
$1-|R|^{-1}+O(|R|^{-2}).$ (19)
For a large solution set, stopping after two iterations can give all solutions with a high probability. However, it is more common to stop after finding no new solutions over a predetermined number of loops. In simulations presented here, we use the monodromy solver in HomotopyContinuation.jl and their default stopping criterion that stops the solver after $5$ loops with no new solutions.